OpenStudy (anonymous):

-5|3z+8|-5=-20 , can anyone help me with the steps on how to check the solutions ?

OpenStudy (anonymous):

steps are 1) add 5 2) divide by -5 3) remove absolute values signs, get two equations 4) solve them

OpenStudy (anonymous):

\[-5|3z+8|-5=-25\] \[-5|3z+8|=-15\] \[|3z+8|=3\]

OpenStudy (anonymous):

those are first two steps. next one is \[3z+8=3\] \[3z=-5\] \[z=-\frac{3}{5}\] or \[3z+8=-3\] \[3z=-11\] \[z=-\frac{11}{3}\]

OpenStudy (anonymous):

check my algebra in solving for x. hope they are clear